Previous | Calendar |
July |
2023 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
25JD176Site Visit (SpB, JvB, BsB, DD): Details , Photos (22) Turned on the fan in the LI-6262 box Nice weather. |
26JD177 |
27JD178Jane S. team setup dust samplers on the tower. |
28JD179 |
29JD180Jane S. team removed dust samplers from the tower. |
30JD181Site Visit (SpB, WW, DL, +): Details , Photos (23) Tour of the US-NR1 site and ecoTram by NCAR CGD. End of the NSF Fluxcourse at MRS. SpB Prius Rain held off till 1pm'ish |
1JD182 |
2JD183 |
3JD184SpB Off (Full "Buck" Moon) |
4JD185Independence Day |
5JD186SpB Off at around 14:39 MST, the 21.5m LI-7500A (op2) and LI-7200 rebooted (likely power outages?) In-cloud/fog (most of day) |
6JD187 |
7JD188SpB Off |
8JD189 |
9JD190 (wet) |
10JD191 |
11JD192 |
12JD193 |
13JD194 |
14JD195Site Visit (SpB): Details , Photos (34) Span cylinder at 500 psi, N2 at 250 psi at 10:45 MST, downloaded PolaRx5 GPS "A" (sn 3062189) data files at 11:30 MST, downloaded PolaRx5 GPS "B" (sn 3227662) data files (in forest) at 11:35 MST, swapped SD card from GPSCAM at 11:20 MST, releveled forest PolaRx5 GPS "B" antenna (see photos) from 11:58-12:05 MST swapped the 21.5m LI-7200 inlet (now using heated inlet) connected AP200 RS232 output to /dev/ttyS12 on isffa (not yet collecting data) NOTE: the 2m REBS Rnet sensor dome is cracked/crushed (need to replace!) Enterprise Leaf Clear in AM; then partly-cloudy |
15JD196 |
16JD197Site Visit (SpB, JvB, BsB, BC): Details , Photos (24) in AM: started ops107; started ingesting the AP-200 data into the DSM Added salt/water to the tower grounding rods PsT BMW Clear skies; a bit of breeze |
17JD198 |
18JD199 |
19JD200The GPS antenna near the ground (PolaRx5 GPS "B") became slightly tilted (see photos) |
20JD201Site Visit (SpB): Details , Photos (16) Span cylinder at 100 ---> 1600 psi, N2 at 100 ---> 1900 psi swapped span; removed ff66921; started cc477126 swapped N2; removed cc254426; started cc428961 started testing of three AM25T loggers in the trailer at C-1 SpB Prius Rain overnight; raining as I left MRS |
21JD202at around 10:55 MST, the 21.5m LI-7500A (op2) and LI-7200 rebooted (likely power outages?) |
22JD203Site Visit by Thai visitors (PB) Span cylinder at 1650 psi, N2 at 1900 psi |
23JD204 |
24JD205SpB Off (Start) |
25JD206 |
26JD207Jane S. team setup dust samplers on the tower. |
27JD208 |
28JD209Jane S. team removed dust samplers from the tower. |
29JD210 |
30JD211 |
31JD212 |
1JD213between 15:10-15:45 MST, the 21.5m LI-7500A (op2) and LI-7200 rebooted (likely power outages?) at around 16:30 MST, the network connection to US-NR1 goes down (Full "Sturgeon" Moon) |
2JD214sometime in AM, the network connection to US-NR1 is back up at around 15:45 MST, the 21.5m LI-7500A (op2) and LI-7200 rebooted (likely power outages?) at around 15:45 MST, the 2.5m CSAT3 stops sending data to the data system (not fixed until Aug XX) |
3JD215 |
4JD216SpB worked (remotely) |
5JD217 |